A random sample of size 4 is taken with replacement from a population with population mean of 12 and population variance of 8. [What is the standard deviation of the sampling distribution]. 


1
Expert's answer
2022-03-24T09:49:47-0400

Standard deviation of the sampling distribution will be "\sigma_x={\frac {\sigma} {\sqrt n}={\frac {\sqrt 8} {\sqrt 4}}}=\sqrt 2"

In the formula above n is sample size and "\sigma" is population standard deviation